# Entanglement between dependent degrees of freedom: Quasi-particle correlations

### v1

@@ III.4 Quantum-correlated eigenstates of non-interacting Hamiltonians
while recalling $\varphi_{1}=\varphi$, $\beta_{1}=\beta$, $N=m+n$, and Eq. (8). For $N>1$, those conditions are generally not met, therefore demonstrating that the eigenstates of an interaction-free Hamiltonian in dependent degrees of freedom can, in fact, be entangled. For example, the case $\delta=0$ applied to the conditions in Eq. (18) simplify them to
@@
which has integer solutions if and only if the total particle number $N$ is a perfect square. Conversely, $\sqrt{N}\notin\mathbb{N}$ leads to entangled $N$-quasi-particle eigenstates, contrary to the separation of eigenstates for independent degrees of freedom without coupling terms in the Hamiltonian.
III.5 Preliminary discussion
We established the notion of classically correlated states for dependent degrees of freedom, being based on products states in Eq. (11). Formally, those states admit the same structure as non-entangled states for independent degrees of freedom, Eq. (10). However, we proved by examples two key distinguishing features of the different kinds of degrees of freedom.
In the first example, we showed linear dependence in the dependent case. This led to the observation of disentangled NOON states, which are always entangled for independent degrees of freedom. The second example pertained to the product nature of eigenstates of non-interacting Hamiltonians, which are products for independent degrees of freedom. Again, a departure from this behavior was observed for dependent degrees of freedom.
@@ III.5 Preliminary discussion
IV Comparing different quantum statistics
Thus far, we compared the separation with respect to independent degrees of freedom, consisting of one fermionic and one bosonic field (i.e., $\mathrm{FB}$ separation), and the quasi-particle description via hybrid eigenmodes (i.e., a separation with respect to the $+$ and -- modes), obeying neither a fermionic nor a bosonic algebra. In this section, we directly compare the previous cases with composite systems consisting purely of fermionic and bosonic quantized fields.
IV.1 Fermions, bosons, and combinations thereof
For the sake of simplicity, we assume a vanishing detuning, $\delta=0$ (likewise, $\omega_{F}=\omega=\omega_{B}$), and a zero phase for the coupling parameter, $\kappa/|\kappa|=e^{i\theta}=1$, throughout this section. For example, this yields balanced superposition [see Eq. (3)] for the quasi-particle modes
@@ IV.1 Fermions, bosons, and combinations thereof
The factorizable $N$-quasi-particle states from Eq. (13) simplify to
@@
Analogously to the previous scenario, Eq. (1), we can consider Hamiltonians $\hat{H}/\hbar=\omega(\hat{f}_{1}^{\dagger}\hat{f}_{1}+\hat{f}_{2}^{\dagger}\hat{f}_{2})+|\kappa|(\hat{f}_{1}^{\dagger}\hat{f}_{2}+\hat{f}_{2}^{\dagger}\hat{f}_{1})/2$ and $\hat{H}/\hbar=\omega(\hat{b}_{1}^{\dagger}\hat{b}_{1}+\hat{b}_{2}^{\dagger}\hat{b}_{2})+|\kappa|(\hat{b}_{1}^{\dagger}\hat{b}_{2}+\hat{b}_{2}^{\dagger}\hat{b}_{1})/2$ for purely fermionic and bosonic systems in second quantization, respectively. The corresponding annihilation and creation operators satisfy the fundamental (anti-)commutation relations $\{\hat{f}_{j},\hat{f}_{j^{\prime}}\}=0=[\hat{b}_{j},\hat{b}_{j^{\prime}}]$ and $\{\hat{f}_{j},\hat{f}_{j^{\prime}}^{\dagger}\}=\delta_{j,j^{\prime}}=[\hat{b}_{j},\hat{b}_{j^{\prime}}^{\dagger}]$ for $j,j^{\prime}\in\{1,2\}$. Furthermore, we find a diagonalization of the Hamiltonians via the eigenmodes
@@
The vacuum states are in all cases identical. Also, the single-particle states, $(|0,1\rangle_{XY}\pm|1,0\rangle_{XY})/\sqrt{2}$, do not show differences for independent modes (i.e., $XY\in\{\mathrm{FF},\mathrm{BB}\}$) and dependent degrees of freedom, $XY=\mathrm{FB}$. Interesting differences appear for $N>1$. For instance, the two-particle states for two-mode fermions are always of a product form, regardless the mode decomposition. Thus, in the following, we continue our comparison of only quasi-particles and bosons for $N=2$.
Figure 3: Comparison of excitations of two bosons and two quasi-particles. The left plot shows the original space spanned by two bosons ($\mathrm{BB}$) and the subspace of the fermion-boson system ($\mathrm{FB}$). The center plot shows the effect of the rotation to the superposition modes, resulting in tensor-product basis states in the purely bosonic case. The right plot shows the projection $\hat{P}$, while preserving normalization, to the constrained quasi-particle states. In the middle plot, all states on the sphere but $|2-n\rangle_{+}\otimes|n\rangle_{-}$ and $-(|2-n\rangle_{+}\otimes|n\rangle_{-})$ for $n\in\{0,1,2\}$ exhibit entanglement of independent degrees of freedom. However, for dependent degrees of freedom, every state on the circle except for the projected points $|2-n,n\rangle_{\mathrm{\pm}}$, together with their reflections, are the entangled ones in the right plot.
IV.2 Two bosons, quasi-particles, and physical projections
@@
For comparison, the explicit expansion of all pure factorizable two-quasi-particle states reads
@@
Up to a normalization, all instances of both cases are only distinguished through the presence ($\mathrm{BB}$ case) and absence ($\mathrm{FB}$ case) of the $|2,0\rangle_{XY}$ component in the original mode decomposition. This makes sense as, for a single fermionic mode, we can identify
@@
with a projection operator in Fock basis as
@@
Note that this relation holds true for $N=m+n>2$, too.
The cases of independent degrees of freedom (two bosonic modes) and dependent degrees of freedom (quasi-particle modes) are related via a projection. This projection also explains the linear dependence since $\hat{P}$ is not bijective and maps an orthonormal tensor-product basis $|m\rangle_{+}\otimes|n\rangle_{-}$ to a linearly dependent set of vectors $|m,n\rangle_{\pm}$. In general, the definition of non-entangled states in Eq. (11) now takes the simpler form
@@
while choosing the scaling of the input states such that the projected result is properly normalized. In this form, the relation of non-entangled states $|x,y\rangle_{\pm}$ to tensor-product states, Eq. (31), becomes clearer. Also, the physical restriction of the composite system is achieved via the projection operator $\hat{P}$, projection onto the fermionic mode $\hat{f}$ as a part of both quasi-particle modes $\hat{p}_{\pm}$. See Fig. 3.
V Generalizations
While we solely considered a specific system of interest so far, we generalize the approach to entanglement between more general dependent degrees of freedom. This includes the treatment of multimode fermion-boson interfaces, a projection-based formalization for other systems, and the mentioning of another application in angular-momentum degrees of freedom.
V.1 Multimode generalization
The framework discussed thus far was based on one fermionic and one bosonic mode. Beyond that, suppose we have $\hat{f}_{j}$, with $j\in\{1,\ldots,M\}$, and $\hat{b}_{j}$ as annihilation operators of our quantized modes in an $2M$-mode composite system. These operators obey the anti-commutator and commutator relations,
@@ V.1 Multimode generalization
Non-entangled states, akin to Eq. (11), now are defined as
@@
A convex hull construction then allows us to express mixed non-entangled (i.e., separable) states as statistical mixtures
@@
Formally, the above definition of multimode non-entangled states for dependent degrees of freedom is identical to that for independent degrees of freedom. Hence, it is consistent with the existing approaches. However, as we have shown in Sec. III for $M=2$, the entanglement and disentanglement of states can deviate from the common expectation.
@@
V.2 Projected product states
As shown in this work, the entanglement characteristics of quasi-particles with dependent degrees of freedom can be determined through non-entangling states that are projected product states, Eq. (30). We can formalize this approach. Let $\mathcal{H}=\hat{P}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$ be a composite Hilbert space subject to a constraint given by the projection $\hat{P}$, with $\hat{P}^{2}=\hat{P}=\hat{P}^{\dagger}$. A pure non-entangled state can be defined as
where $|x\rangle_{X}\in\mathcal{H}_{X}$ and $|y\rangle_{Y}\in\mathcal{H}_{Y}$ are local states of the joint system. Note that this definition is identical with the one for a factorizable state for independent degrees of freedom, where $\hat{P}=\hat{1}$. An entangled state $|\Psi\rangle$ does not admit the decomposition from Eq. (36), $|\Psi\rangle\neq\hat{P}\,(|x\rangle_{X}\otimes|y\rangle_{Y})$.
Furthermore, a mixed, classically correlated state reads
[…]
And a state that does not obey this form is consequently inseparable, i.e., entangled. Also note that this definition straightforwardly generalizes to multipartite systems. Therefore, a separable state is defined as a state that obeys the physical constraint (given by the projector $\hat{P}$) and only contains classical correlations (via the joint probability distribution $P$) of local tensor-product states $|x\rangle_{X}\otimes|y\rangle_{Y}$.
In addition to the quasi-particle description, whose projection can be found in Eq. (29), another example is in order. Consider the degree of freedom pertaining to the total angular momentum, $J^{2}$, and one for the angular momentum in $z$ direction, $J_{z}$. It is well known that the corresponding angular-momentum operators can be jointly diagonalized,
$\text{with}\quad-j\leq m\leq j$ $\quad\text{and}\quad j\geq 0.$
where $j$ and $m$ could, in principle, take arbitrary (half) integer values without the restrictions. The projection $\hat{P}$ then makes sure that the angular momentum in one direction never exceeds the total angular momentum, $|m|\ngtr j$. The above definition further allows us to define non-entangled states,
As a finial remark, it is worth mentioning that $\hat{P}$ does not have to be a projection operator for our above general definitions of separability for dependent degrees of freedom to work. Analogously to the aforementioned scenario of partially overlapping modes, we can generalize $\hat{P}$ to arbitrary linear operators if required by a physical constraint that yields the composite system, $\mathcal{H}=\hat{P}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$.
VI Conclusion
In this work, we explored the impact of dependent degrees of freedom on quantum entanglement properties. Defying common expectations, we showed that certain states are non-entangled in the case of dependent degrees of freedom which are entangled for independent notions of entanglement, and vice versa.
The main physical system to which we applied our approach were quasi-particle descriptions for light-matter quantum interfaces via a Jaynes–Cummings model. We showed that non-entangled quasi-particle states can be expressed through projections—accounting for physical constraints—that act on tensor-product states. The projection leads to linear dependencies, enabling us, for example, to construct NOON-type states which are not entangled, unlike their counterparts for independent degrees of freedom. Conversely, the true statement that eigenvectors of a local (and non-degenerate) Hamiltonian are necessarily factorizable is generally not true for dependent degrees of freedom.
Furthermore, we compared our fermion-boson-composite quasi-particle description with purely bosonic and fermionic systems. For a single excitation, no distinction could be made between the different cases. However, we were able to identify unique entanglement characteristics for the different physical scenarios when considering at least two (quasi-)particles. Eventually, a generalization to multipartite entanglement of dependent degrees of freedom, including the treatment of mixed states and physical systems other than quasi-particles, were established.
@@ Acknowledgements.
The authors acknowledge funding through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078). This work was further supported through the Ministerium für Kultur und Wissenschaft des Landes Nordrhein-Westfalen through the project PhoQC: Photonisches Quantencomputing.

### v2

@@ IV.4 Quantum-correlated eigenstates of noninteracting Hamiltonians
while recalling $\varphi_{1}=\varphi$, $\beta_{1}=\beta$, $N=m+n$, and Eq. (13). For $N>1$, those conditions are generally not met, therefore demonstrating that the eigenstates of an interaction-free Hamiltonian in dependent degrees of freedom can, in fact, be entangled. For example, the case $\delta=0$ applied to Eq. (23) simplifies the conditions therein to
@@
which have integer solutions if and only if the total particle number $N$ is a perfect square. Conversely, $\sqrt{N}\notin\mathbb{N}$ leads to entangled $N$-quasiparticle eigenstates, contrary to the separation of eigenstates for independent degrees of freedom without coupling terms in the Hamiltonian.
IV.5 Discussion
We established the notion of classically correlated states for dependent degrees of freedom, being based on products states in Eq. (16). Formally, those states admit the same structure as nonentangled states for independent degrees of freedom [Eq. (15)]. However, we proved by examples two key distinguishing features of the different kinds of degrees of freedom.
In the first example, we showed linear dependence in the dependent case. This led to the observation of disentangled NOON states, which are always entangled for independent degrees of freedom. The second example pertained to the product nature of eigenstates of noninteracting Hamiltonians, which are products for independent degrees of freedom. Again, a departure from this behavior was observed for dependent degrees of freedom.
@@ IV.5 Discussion
V Comparing different quantum statistics
Thus far, we compared the separation with respect to independent degrees of freedom, consisting of one fermionic and one bosonic field (i.e., $\mathrm{FB}$ separation), and the quasiparticle description via hybrid eigenmodes (i.e., a separation with respect to the $+$ and -- modes), obeying neither a fermionic nor a bosonic algebra. In this section, we directly compare the previous cases with composite systems consisting purely of fermionic and bosonic quantized fields.
V.1 Fermions, bosons, and combinations thereof
For the sake of simplicity, we assume a vanishing detuning, $\delta=0$ (likewise, $\omega_{F}=\omega=\omega_{B}$), and a zero phase for the coupling parameter, $\kappa/|\kappa|=e^{i\theta}=1$, throughout this section. For example, this yields balanced superposition [see Eq. (8)] for the quasiparticle modes
@@ V.1 Fermions, bosons, and combinations thereof
The factorizable $N$-quasiparticle states from Eq. (18) simplify to
@@
Analogously to the previous scenario, Eq. (6), we can consider Hamiltonians $\hat{H}/\hbar=\omega(\hat{f}_{1}^{\dagger}\hat{f}_{1}+\hat{f}_{2}^{\dagger}\hat{f}_{2})+|\kappa|(\hat{f}_{1}^{\dagger}\hat{f}_{2}+\hat{f}_{2}^{\dagger}\hat{f}_{1})/2$ and $\hat{H}/\hbar=\omega(\hat{b}_{1}^{\dagger}\hat{b}_{1}+\hat{b}_{2}^{\dagger}\hat{b}_{2})+|\kappa|(\hat{b}_{1}^{\dagger}\hat{b}_{2}+\hat{b}_{2}^{\dagger}\hat{b}_{1})/2$ for purely fermionic and bosonic systems in second quantization, respectively. The corresponding annihilation and creation operators satisfy the fundamental (anti)commutation relations $\{\hat{f}_{j},\hat{f}_{j^{\prime}}\}=0=[\hat{b}_{j},\hat{b}_{j^{\prime}}]$ and $\{\hat{f}_{j},\hat{f}_{j^{\prime}}^{\dagger}\}=\delta_{j,j^{\prime}}=[\hat{b}_{j},\hat{b}_{j^{\prime}}^{\dagger}]$ for $j,j^{\prime}\in\{1,2\}$. Furthermore, we find a diagonalization of the Hamiltonians via the eigenmodes
@@
The vacuum states are in all cases identical. Also, the single-particle states, $(|0,1\rangle_{XY}\pm|1,0\rangle_{XY})/\sqrt{2}$, do not show differences for independent modes (i.e., $XY\in\{\mathrm{FF},\mathrm{BB}\}$) and dependent degrees of freedom, $XY=\mathrm{FB}$. Interesting differences appear for $N>1$. For instance, the two-particle states for two-mode fermions are always of a product form, regardless of the mode decomposition. Thus, in the following, we continue our comparison of only quasiparticles and bosons for $N=2$.
Figure 3: Comparison of excitations of two bosons and two quasiparticles. The left plot shows the original space spanned by two bosons ($\mathrm{BB}$) and the subspace of the fermion-boson system ($\mathrm{FB}$). The center plot shows the effect of the rotation to the superposition modes, resulting in tensor-product basis states in the purely bosonic case. The right plot shows the projection $\hat{P}$, while preserving normalization, to the constrained quasiparticle states. In the middle plot, all states on the sphere but $|2-n\rangle_{+}\otimes|n\rangle_{-}$ and $-(|2-n\rangle_{+}\otimes|n\rangle_{-})$ for $n\in\{0,1,2\}$ exhibit entanglement of independent degrees of freedom. However, for dependent degrees of freedom, every state on the circle except for the projected points $|2-n,n\rangle_{\mathrm{\pm}}$, together with their reflections, are the entangled ones in the right plot.
V.2 Two bosons, quasiparticles, and physical projections
@@
For comparison, the explicit expansion of all pure factorizable two-quasiparticle states reads
@@
Up to a normalization, all instances of both cases are only distinguished through the presence ($\mathrm{BB}$ case) and the absence ($\mathrm{FB}$ case) of the $|2,0\rangle_{XY}$ component in the original mode decomposition. This makes sense as, for a single fermionic mode, we can identify
@@
with a projection operator in the Fock basis as
@@
Note that this relation holds true for $N=m+n>2$, too. Eventually, this presents the sought relation to the notion of dependent degrees of freedom, established in Sec. II.
The cases of independent degrees of freedom (two bosonic modes) and dependent degrees of freedom (quasiparticle modes) are related via a projection. This projection also explains the linear dependence since $\hat{P}$ is not bijective and maps an orthonormal tensor-product basis $|m\rangle_{+}\otimes|n\rangle_{-}$ to a linearly dependent set of vectors $|m,n\rangle_{\pm}$. In general, the definition of nonentangled states in Eq. (16) now takes the simpler form
@@
while choosing the scaling of the input states such that the projected result is properly normalized. In this form, the relation of nonentangled states $|x,y\rangle_{\pm}$ to tensor-product states, Eq. (36), becomes clearer. Also, the physical restriction of the composite system is achieved via the projection operator $\hat{P}$, projection onto the fermionic mode $\hat{f}$ as a part of both quasiparticle modes $\hat{p}_{\pm}$. See Fig. 3 for a visualization.
VI Multimode generalization
The framework discussed thus far is based on one fermionic and one bosonic mode. Beyond that, suppose we have $\hat{f}_{j}$, with $j\in\{1,\ldots,M\}$, and $\hat{b}_{j}$ as annihilation operators of our quantized modes in a $2M$-mode composite system. These operators obey the anticommutator and commutator relations,
@@ VI Multimode generalization
Nonentangled states, akin to Eq. (16), now are defined as
@@
A convex hull construction then allows us to express mixed nonentangled (i.e., separable) states as statistical mixtures
@@
Formally, the above definition of multimode nonentangled states for dependent degrees of freedom is identical to that for independent degrees of freedom. Hence, it is consistent with the existing approaches. However, as we have shown in Sec. IV for $M=2$, the entanglement and disentanglement of states can deviate from the common expectation.
@@
VII Conclusion
In summary, we established and explored entanglement for dependent degrees of freedom, which is compared to the archetypal notion of entanglement between independent degrees of freedom. Defying common expectations, we showed that certain states are nonentangled in the case of dependent degrees of freedom which are entangled for independent notions of entanglement, and vice versa.
The main physical systems to which we applied our approach were quasiparticle descriptions for light-matter quantum interfaces via a Jaynes–Cummings model. We showed that nonentangled quasiparticle states can be expressed through projections—accounting for physical constraints—that act on tensor-product states. The projection leads to linear dependencies, enabling us, for example, to construct NOON-type states which are not entangled, unlike their counterparts for independent degrees of freedom. Conversely, the true statement that eigenvectors of a local (and nondegenerate) Hamiltonian are necessarily factorizable is generally not true for dependent degrees of freedom.
Furthermore, we compared our fermion-boson-composite quasiparticle description with purely bosonic and fermionic systems. For a single excitation, no distinction could be made between the different cases. However, we were able to identify unique entanglement characteristics for the different physical scenarios when considering at least two (quasi)particles. Eventually, a generalization to multipartite entanglement of dependent degrees of freedom, including the treatment of mixed states and physical systems other than quasiparticles, was established.
@@ Acknowledgements.
The authors acknowledge funding through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Transregional Collaborative Research Center TRR 142 (Projects No. A04 and No. C10, Grant No. 231447078). This work was further supported through the Ministerium für Kultur und Wissenschaft des Landes Nordrhein-Westfalen through the project PhoQC: Photonisches Quantencomputing.